40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Wed Aug 5 19:03:56 2020, aaron, Update, Lab Work, Notice of lab entry 
    Reply  Fri Aug 7 11:35:09 2020, aaron, Update, Lab Work, lab work 
Message ID: 2551     Entry time: Fri Aug 7 11:35:09 2020     In reply to: 2550
Author: aaron 
Type: Update 
Category: Lab Work 
Subject: lab work 

Entered the lab at 11:00 am. The lab is far too hot (78F) and humid (45%)! 

Inventoried available photodiodes in the cryo lab, on the PSOMA wiki.

Fri Aug 7 12:55:53 2020
Facilities came to check the fire alarm. I let them in with shoe covers, they held a magnet up to the fire alarm and left. Total operation took less than 5 min.
[lunch and a couple meetings 1 - 3]

Pretty tired honestly, I submitted an order for a few things we want in the lab, such as:

  • 2x "ultra quiet" fans to circulate air
  • longer ethernet cable for reaching far end of PSOMA table
  • headset web cam so we can take video while doing lab activities
  • more plastic storage bins
  • food and water for our emergency kit

The PSOMA optics table is from TMC vibration control (TMC 784-29701-01). I sent them an inquiry for the datasheet, and also asked whether they sell clean enclosures or have application notes that might help us put one together. Also talked with Arian Jadbabaie a bit about the Hutzler lab enclosures, and got some photos.

Exit ~430 pm

ELOG V3.1.3-